OpenStudy (anonymous):

A triangle has a base of 13 meters and a height of 6 meters. What is the area of the triangle? A. 39 square meters B. 169 square meters C. 156 square meters D. 78 square meters

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }*base*height\]

OpenStudy (anonymous):

So \[\frac{ 1 }{ 2 }*13*6=answer\]

sammixboo (sammixboo):

Right and when multiplying something by 1/2 you just take the number and half it, so 1/2*6 = 3 and 1/2*5 = 2.5